write an equation for. Mrs Smith rode the bus 720 miles in 3 days. On the first day she rode 3 times as far as she did on the second day. On the third day, she rode 2 times as far as she did the second day. How far did she travel each day?
Answer by haileytucki(390)

x+3x+2x=720
Since x and 3x are like terms, add 3x to x to get 4x.
4x+2x=720
Since 4x and 2x are like terms, add 2x to 4x to get 6x.
6x=720
Divide each term in the equation by 6.
(6x)/(6)=(720)/(6)
Simplify the left-hand side of the equation by canceling the common factors.
x=(720)/(6)
Simplify the right-hand side of the equation by simplifying each term.
x=120
First day 120*3=360 miles
Second Day = 120 miles
Thrid Day=120*2=240 miles